OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 A. {0.4, 5.6} B. {0.5, 4.8} C. {1.4, 2.4} D. no real number solution Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 A. {0.4, 5.6} B. {0.5, 4.8} C. {1.4, 2.4} D. no real number solution @Mathematics

OpenStudy (amistre64):

you can plug in the first numbers and test it out to see whats closest

OpenStudy (amistre64):

36-8 aint negative so it has real solutions

OpenStudy (amistre64):

.4+5.6 = 6/2 = 3 ; and 6/2 = 3 but thats might be a lucky shot .5+4.8 = 5.3/2 is less than 3 hmmm and 1.2 + 2.4 aint got 3 inbetween them. your axis of symmetry is 3 so id go a

OpenStudy (amistre64):

the wolf agrees http://www.wolframalpha.com/input/?i=x2+-+6x+%2B+2+%3D+0